# Algorithms for Approximating Conditionally Optimal Bounds

### v3

@@ 1 Preliminaries
A partial order is a preorder where $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\sim_{R}{\bm{y}}$ implies that ${\bm{x}}={\bm{y}}$. When $\bm{x}$} is no greater than $\bm{y}$} in partial order $R$ we write ${\bm{x}}\leq_{R}{\bm{y}}$. We further user ${\bm{x}}<_{R}{\bm{y}}$ to indicate that ${\bm{x}}\leq_{R}{\bm{y}}$ and ${\bm{x}}\neq{\bm{y}}$.
@@
A total order is a partial order that is strongly connected, i.e., $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\leq_{T}{\bm{y}}$ or ${\bm{y}}\leq_{T}{\bm{x}}$.
@@
Let $R$ be an arbitrary preorder and take any ${\bm{x}},{\bm{y}}\in\Omega$. We have
${\bm{x}}<_{R}{\bm{y}}\Leftrightarrow\neg({\bm{y}}\lesssim_{R}{\bm{x}}).$ (2)
@@
We prove the contrapositive of the reverse direction: ${\bm{y}}\lesssim_{R}{\bm{x}}\Rightarrow\neg({\bm{x}}<_{R}{\bm{y}})$. Suppose again for the sake of contradiction that ${\bm{y}}\lesssim_{R}{\bm{x}}$ and ${\bm{x}}<_{R}{\bm{y}}$. The latter implies that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$. But together ${\bm{y}}\lesssim_{R}{\bm{x}}$ and ${\bm{x}}\lesssim_{R}{\bm{y}}$ imply by Definition 1 that ${\bm{x}}\not\sim_{R}{\bm{y}}$, a contradiction. ∎
@@
The upper set associated with sample $\bm{x}\in\Omega$ and preorder $R$ is given by
@@
Note that, according to these definitions, $\forall\bm{x},\bm{y}\in\Omega$, preorder $R$, and total order $T$, $\bm{x}\lesssim_{R}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},R)$ and ${\bm{x}}\leq_{T}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},T)$.
@@ 1.1 Common orders
The low lexicographic order, $T_{\ell}$, is such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{\ell}}\bm{y}\Leftrightarrow L_{1}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
@@
The high lexicographic order, $T_{h}$, is such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{h}}\bm{y}\Leftrightarrow H_{n}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
@@
For fixed $i\in\{1,\ldots,n\}$ and any two samples ${\bm{x}},{\bm{y}}\in\Omega$, the $i$th quantile preorder $R_{i}$ is defined by
@@
Let $R$ be an arbitrary weakly monotone partial order. By Definitions 2 and 5 we have that
@@
All orders discussed in this document are strongly monotone. For brevity, we refer to such orders as simply monotone going forward.
@@ 2 Topology
Take any $F\in\mathcal{F}(\Omega^{\prime},\alpha)$. By the definition of closure we know that $\forall\delta>0,\exists G\in\mathcal{G}(\Omega^{\prime},\alpha):d(G,F)<\delta$. Choose $\delta=\epsilon/(\sqrt{m}S_{\max})$. According to Inequality 12, we have $|E[F]-E[G]|<\epsilon$. ∎
@@
Suppose first that there exists $F\in\mathcal{F}(\Omega^{\prime},\alpha)$ such that $E[F]=\mu$ and choose any $\epsilon>0$. By the definition of closure we know that $\forall i>0,\exists G_{i}\in\mathcal{G}(\Omega^{\prime},\alpha):d(G_{i},F)<1/i$. According to Inequality 12, when $i>\sqrt{m}S_{\max}/\epsilon$, $|E[F]-E[G_{i}]|<\epsilon$. Thus, allowing $\epsilon$ to approach 0, we can see that the sequence of distributions $\{G_{i}\}_{i>0}$ converges in mean to $F$.
@@ 3 Conditional-Optimality
If total order $T$ agrees with preorder $R$, then $\forall{\bm{x}}\in\Omega$, $\Omega({\bm{x}},T)\subseteq\Omega({\bm{x}},R)$.
@@
Let ${\bm{x}}\in\Omega$ be arbitrary and note that by Definition 12, if total order $T$ agrees with preorder $R$, then for each ${\bm{y}}\in\Omega$
@@
For every preorder $R$, there exists a total order $T$ that agrees with $R$.
Proof.
Every preorder $R$ is a partial order over equivalence classes. Construct partial order $D$ by maintaining all non-equivalence relations of $R$ and arbitrarily assigning a total order between the elements within each of the equivalence classes of $R$. Now take any ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{R}{\bm{y}}$. Since ${\bm{x}}\not\sim_{R}{\bm{y}}$, it’s clear by construction that ${\bm{x}}<_{D}{\bm{y}}$ as well. Thus, partial order $D$ agrees with $R$.
Proposition 7.
For every total order $T$ agreeing with preorder $R$ and any ${\bm{x}}\in\Omega$, there exists a ${\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\Omega({\bm{x}},R)=\Omega({\bm{y}},T)$.
@@
Bound $B$ is consistent with preorder $R$ if and only if $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ and $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$.
@@
Now suppose that $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$ and $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ for preorder $R$. Let $T$ be any total order agreeing with $R$. For each ${\bm{x}},{\bm{y}}\in\Omega$ we have by Proposition 1 and taking the contrapositive of Definition 12 that
@@
If ${\bm{x}}<_{R}{\bm{y}}$, then by assumption $B({\bm{x}})\leq B({\bm{y}})$. And if ${\bm{x}}\sim_{R}{\bm{y}}$, we again have by assumption that $B({\bm{x}})=B({\bm{y}})\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Thus, in any case, ${\bm{x}}\lesssim_{R}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. It follows then by Implication 16 that ${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Therefore, by Definition 13, it must be the case that $B$ is consistent with $T$. Since $T$ was an arbitrary total order agreeing with $R$, we have, by Definition 14, that $B$ is also consistent with $R$. ∎
@@
For any preorder $R$, $B_{R}^{*}$ is consistent with $R$.
@@
Lemma 5.
$B_{R}^{*}$ is the conditionally optimal bound for preorder $R$ provided that $\mathcal{F}(\Omega({\bm{x}},R))$ is non-empty.
@@
We begin by noting that, according to Lemma 4, $B_{R}^{*}$ is consistent with $R$. Thus, by Definition 15, it suffices to show that for any valid bound $B_{R}$, consistent with $R$, and $\forall{\bm{x}}\in\Omega,B^{*}_{R}({\bm{x}})\geq B_{R}({\bm{x}})$. This would imply that $B_{R}$ is either strictly worse than $B_{R}^{*}$ for some ${\bm{x}}\in\Omega$, or that $B_{R}=B_{R}^{*}$. To that end, suppose to the contrary that there exists some valid $B_{R}$, also consistent with $R$, such that $\exists{\bm{x}}\in\Omega:B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. Now let $T$ be any total order agreeing with $R$ and take ${\bm{y}}\in\Omega$ such that ${\bm{y}}\sim_{R}{\bm{x}}$ and $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)=\Omega({\bm{x}},R)$. We are guaranteed that such a $T$ exists by Propositions 6, and we are guaranteed that such a $\bm{y}$} exists by Proposition 7. Since ${\bm{x}}\sim_{R}{\bm{y}}$ and $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$, it’s clear by Lemma 3 that
@@
And because $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)$ we also have by Definition 16 that $B_{T}^{*}({\bm{y}})=B_{R}^{*}({\bm{y}})$. So it must be the case that $B_{T}^{*}({\bm{x}})<B_{R}({\bm{y}})$ as well, which is a contradiction given that $B_{T}^{*}$ is conditionally optimal by Theorem 1. ∎
@@ 5 Monotone Orders
For any monotone order $T$ and homogeneous sample $\bm{S}_{i}\in\Omega$ we have
@@
Let $T$ be any monotone order, $B_{T}^{*}$ the conditionally optimal bound for $T$, and $\bm{x}\in\Omega$ an arbitrary sample. If $S_{i}\in S$ is such that $\bm{S}_{i}\leq_{T}\bm{x}\leq_{T}\bm{S}_{i+1}$, then
@@
For any homogeneous sample $S_{i}\in S$, $B_{T_{h}}^{*}(\bm{S}_{i})$ is the weakest (lowest) and $B_{T_{\ell}}^{*}(\bm{S}_{i})$ the strongest (highest) among all bounds conditionally optimal with respect to a monotone order.
@@ 5.1 Calculating bounds for homogeneous samples
From Learned-Miller [2] Lemma 2.2, the result is true for $S_{i}=S_{\min}$, so we fix $\bm{S}_{i}\neq\bm{S}_{\min}$. By Theorem 1 we need only show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}=\mu$ falls within the interval defined by Inequality 32. To that end, by Proposition 9, it will suffice to show the following two properties. On one hand, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}$, $E[H]>S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})$. And on the other hand, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose limiting mean is bounded above by $S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha})$.
@@ 6 Refinements of $\mathcal{F}$
Fix ${\bm{x}}\in\Omega$, $\alpha\in[0,1)$, and preorder $R$. Suppose that there exists a set $C\subseteq S$ such that for any $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},R)]=P_{H}[\Omega({\bm{x}},R)]$. Then we have that
@@
Also by construction and Definition 9, $P_{G}[x_{(i)}\leq X_{(i)}]=P_{H_{p}}[x_{(i)}\leq X_{(i)}]$. Therefore, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$. Take $H_{p^{*}}$ to be the distribution in $\mathcal{F}_{C^{-}}$ corresponding to $G^{*}$. By letting $\varepsilon\rightarrow 0$ and by Inequality 39 and the definition of $G^{*}$ we have

### v4

@@ 1 Preliminaries
Partial order $R$ is a preorder where $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\sim_{R}{\bm{y}}$ implies that ${\bm{x}}={\bm{y}}$. When $\bm{x}$} is no greater than $\bm{y}$} in partial order $R$ we write ${\bm{x}}\leq_{R}{\bm{y}}$. We further user ${\bm{x}}<_{R}{\bm{y}}$ to indicate that ${\bm{x}}\leq_{R}{\bm{y}}$ and ${\bm{x}}\neq{\bm{y}}$.
@@
Total order $T$ is a partial order that is strongly connected, i.e., $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\leq_{T}{\bm{y}}$ or ${\bm{y}}\leq_{T}{\bm{x}}$.
Definition 4.
Total preorder $R$ is a preorder that is strongly connected, i.e. where $\forall{\bm{x}},{\bm{y}}\in\Omega,{\bm{x}}\lesssim_{R}{\bm{y}}$ or ${\bm{y}}\lesssim_{R}{\bm{x}}$. A total preorder is a total order over equivalence classes.
@@
If $R$ is a preorder, then for any ${\bm{x}},{\bm{y}}\in\Omega$. We have
${\bm{x}}<_{R}{\bm{y}}\Rightarrow\neg({\bm{y}}\lesssim_{R}{\bm{x}}).$ (2)
If additionally $R$ is a total preorder, then we also have
$\neg({\bm{y}}\lesssim_{R}{\bm{x}})\Rightarrow{\bm{x}}<_{R}{\bm{y}}.$ (3)
@@
To prove the reverse direction, suppose that $R$ is a total preorder and $\neg({\bm{y}}\lesssim_{R}{\bm{x}})$. Since $R$ is a total preorder, we know by Definition 4 that either ${\bm{x}}\lesssim_{R}{\bm{y}}$ or ${\bm{y}}\lesssim_{R}{\bm{x}}$. Therefore, it must be the case that ${\bm{x}}\lesssim_{R}{\bm{y}}$. Now Definition 1 requires that ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{y}}\lesssim_{T}{\bm{x}}$ in order for ${\bm{y}}\sim_{R}{\bm{x}}$. Therefore ${\bm{x}}\not\sim_{R}{\bm{y}}$, since $\neg({\bm{y}}\lesssim_{R}{\bm{x}})$. Thus, ${\bm{x}}\lesssim_{R}{\bm{y}}$ and ${\bm{x}}\not\sim_{R}{\bm{y}}$ so that ${\bm{x}}<_{R}{\bm{y}}$ by Definition 1. ∎
@@
The upper set associated with sample $\bm{x}\in\Omega$ and total preorder $R$ is given by
@@
Note that, according to these definitions, $\forall\bm{x},\bm{y}\in\Omega$, total preorder $R$, and total order $T$, $\bm{x}\lesssim_{R}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},R)$ and ${\bm{x}}\leq_{T}\bm{y}\Leftrightarrow\bm{y}\in\Omega(\bm{x},T)$.
@@ 1.1 Common orders
The low lexicographic order, $T_{\ell}$, is the total order such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{\ell}}\bm{y}\Leftrightarrow L_{1}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
@@
The high lexicographic order, $T_{h}$, is the total order such that $\forall\bm{x},\bm{y}\in\Omega,\bm{x}\leq_{T_{h}}\bm{y}\Leftrightarrow H_{n}(\bm{x},\bm{y})=1$ where $\forall i\in[n]$,
@@
For fixed $i\in\{1,\ldots,n\}$ and any two samples ${\bm{x}},{\bm{y}}\in\Omega$, the $i$th quantile preorder $R_{i}$ is the total preorder defined by
@@
Let $R$ be an arbitrary weakly monotone partial order. By Definitions 2 and 6 we have that
@@
All monotone orders discussed in this document are strongly monotone. For brevity, we refer to such orders as simply monotone going forward.
@@ 2 Topology
Take any $F\in\mathcal{F}(\Omega^{\prime},\alpha)$. By the definition of closure we know that $\forall\delta>0,\exists G\in\mathcal{G}(\Omega^{\prime},\alpha):d(G,F)<\delta$. Choose $\delta=\epsilon/(\sqrt{m}S_{\max})$. According to Inequality 13, we have $|E[F]-E[G]|<\epsilon$. ∎
@@
Suppose first that there exists $F\in\mathcal{F}(\Omega^{\prime},\alpha)$ such that $E[F]=\mu$ and choose any $\epsilon>0$. By the definition of closure we know that $\forall i>0,\exists G_{i}\in\mathcal{G}(\Omega^{\prime},\alpha):d(G_{i},F)<1/i$. According to Inequality 13, when $i>\sqrt{m}S_{\max}/\epsilon$, $|E[F]-E[G_{i}]|<\epsilon$. Thus, allowing $\epsilon$ to approach 0, we can see that the sequence of distributions $\{G_{i}\}_{i>0}$ converges in mean to $F$.
@@ 3 Conditional-Optimality
For every preorder $R$, there exists a total order $T$ that agrees with $R$.
@@
Every preorder $R$ is a partial order over equivalence classes. Construct partial order $D$ by maintaining all non-equivalence relations of $R$ and arbitrarily assigning a total order between the elements within each of the equivalence classes of $R$ (that such a total order exists follows from the well-ordering principle). Now take any ${\bm{x}},{\bm{y}}\in\Omega$ such that ${\bm{x}}<_{R}{\bm{y}}$. Since ${\bm{x}}\not\sim_{R}{\bm{y}}$, it’s clear by construction that ${\bm{x}}<_{D}{\bm{y}}$ as well. Thus, partial order $D$ agrees with $R$.
Proposition 6.
If total order $T$ agrees with total preorder $R$, then $\forall{\bm{x}}\in\Omega$, $\Omega({\bm{x}},T)\subseteq\Omega({\bm{x}},R)$.
Proof.
Let ${\bm{x}}\in\Omega$ be arbitrary and note that by Definition 13, if total order $T$ agrees with total preorder $R$, then for each ${\bm{y}}\in\Omega$
@@
For every total order $T$ agreeing with total preorder $R$ and any ${\bm{x}}\in\Omega$, there exists a ${\bm{y}}\in\Omega$ such that ${\bm{x}}\sim_{R}{\bm{y}}$ and $\Omega({\bm{x}},R)=\Omega({\bm{y}},T)$.
@@
If bound $B$ is consistent with preorder $R$ then for any ${\bm{x}},{\bm{y}}\in\Omega$, $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ and $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$. The converse also holds if $R$ is a total preorder.
@@
Now suppose that $R$ is a total preorder, $B({\bm{x}})=B({\bm{y}})$ whenever ${\bm{x}}\sim_{R}{\bm{y}}$, and $B({\bm{x}})\leq B({\bm{y}})$ whenever ${\bm{x}}<_{R}{\bm{y}}$ for preorder $R$. Let $T$ be any total order agreeing with $R$. For each ${\bm{x}},{\bm{y}}\in\Omega$ we have by Proposition 1 and taking the contrapositive of Definition 13 that
@@
If ${\bm{x}}<_{R}{\bm{y}}$, then by assumption $B({\bm{x}})\leq B({\bm{y}})$. And if ${\bm{x}}\sim_{R}{\bm{y}}$, we again have by assumption that $B({\bm{x}})=B({\bm{y}})\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Thus, in any case, ${\bm{x}}\lesssim_{R}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. It follows then by Implication 17 that ${\bm{x}}\leq_{T}{\bm{y}}\Rightarrow B({\bm{x}})\leq B({\bm{y}})$. Therefore, by Definition 14, it must be the case that $B$ is consistent with $T$. Since $T$ was an arbitrary total order agreeing with $R$, we have, by Definition 15, that $B$ is also consistent with $R$. ∎
@@
For any total preorder $R$, $B_{R}^{*}$ is consistent with $R$.
@@
Theorem 2.
For every $\alpha\in[0,1)$ and total preorder $R$, $B_{R}^{*}$ is the conditionally optimal bound for $R$ provided that $\mathcal{F}(\Omega({\bm{x}},R),\alpha)$ is non-empty.
@@
We begin by noting that, according to Lemma 4, $B_{R}^{*}$ is consistent with $R$. Thus, by Definition 16, it suffices to show that for any valid bound $B_{R}$, consistent with $R$, and $\forall{\bm{x}}\in\Omega,B^{*}_{R}({\bm{x}})\geq B_{R}({\bm{x}})$. This would imply that $B_{R}$ is either strictly worse than $B_{R}^{*}$ for some ${\bm{x}}\in\Omega$, or that $B_{R}=B_{R}^{*}$. To that end, suppose to the contrary that there exists some valid $B_{R}$, also consistent with $R$, such that $\exists{\bm{x}}\in\Omega:B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. Now let $T$ be any total order agreeing with $R$ and take ${\bm{y}}\in\Omega$ such that ${\bm{y}}\sim_{R}{\bm{x}}$ and $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)=\Omega({\bm{x}},R)$. We are guaranteed that such a $T$ exists by Propositions 5, and we are guaranteed that such a $\bm{y}$} exists by Proposition 7. Since $T$ agrees with $R$ we know by Definition 15 that $B_{R}$ is consistent with $T$. Notice also that Theorem 1 ensures that $B_{T}^{*}$ is conditionally optimal for $T$.
Now since ${\bm{x}}\sim_{R}{\bm{y}}$ and $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$, it’s clear by Lemma 3 that
@@
And because $\Omega({\bm{y}},T)=\Omega({\bm{y}},R)$ we also have by Definition 17 that $B_{T}^{*}({\bm{y}})=B_{R}^{*}({\bm{y}})$. So it must be the case that $B_{T}^{*}({\bm{y}})<B_{R}({\bm{y}})$ as well, which means that there is a bound consistent with $T$ (the bound $B_{R}$) that for sample $\bm{y}$} improves on the bound given by $B_{T}^{*}$. But this contradicts the fact that $B_{T}^{*}$ is conditionally optimal for order $T$. Thus, it cannot be the case that $B_{R}^{*}({\bm{x}})<B_{R}({\bm{x}})$. ∎
@@ 5 Monotone Orders
For any monotone total order $T$ and homogeneous sample $\bm{S}_{i}\in\Omega$ we have
@@
Let $T$ be any monotone total order, $B_{T}^{*}$ the conditionally optimal bound for $T$, and $\bm{x}\in\Omega$ an arbitrary sample. If $S_{i}\in S$ is such that $\bm{S}_{i}\leq_{T}\bm{x}\leq_{T}\bm{S}_{i+1}$, then
@@
For any homogeneous sample $S_{i}\in S$, $B_{T_{h}}^{*}(\bm{S}_{i})$ is the weakest (lowest) and $B_{T_{\ell}}^{*}(\bm{S}_{i})$ the strongest (highest) among all bounds conditionally optimal with respect to a monotone total order.
@@ 5.1 Calculating bounds for homogeneous samples
From Learned-Miller [2] Lemma 2.2, the result is true for $S_{i}=S_{\min}$, so we fix $\bm{S}_{i}\neq\bm{S}_{\min}$. By Theorem 1 we need only show that $\min\{E[F]:F\in\mathcal{F}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}=\mu$ falls within the interval defined by Inequality 33. To that end, by Proposition 9, it will suffice to show the following two properties. On one hand, $\forall H\in\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)\}$, $E[H]>S_{\min}\sqrt[n]{1-\alpha}+S_{i}(1-\sqrt[n]{1-\alpha})$. And on the other hand, there exists a sequence of distributions in $\mathcal{G}(\Omega(\bm{S}_{i},T_{h}),\alpha)$ whose limiting mean is bounded above by $S_{\min}\sqrt[n]{1-\alpha}+S_{i+1}(1-\sqrt[n]{1-\alpha})$.
@@ 6 Refinements of $\mathcal{F}$
Fix ${\bm{x}}\in\Omega$, $\alpha\in[0,1)$, and total preorder $R$. Suppose that there exists a set $C\subseteq S$ such that for any $G,H\in\mathcal{F}$ agreeing cumulatively and pointwise on $C$, $P_{G}[\Omega({\bm{x}},R)]=P_{H}[\Omega({\bm{x}},R)]$. Then we have that
@@
Also by construction and Definition 10, $P_{G}[x_{(i)}\leq X_{(i)}]=P_{H_{p}}[x_{(i)}\leq X_{(i)}]$. Therefore, $H_{p}\in\mathcal{G}_{C^{+}}(\Omega({\bm{x}},R_{i}),\alpha)$. Take $H_{p^{*}}$ to be the distribution in $\mathcal{F}_{C^{-}}$ corresponding to $G^{*}$. By letting $\varepsilon\rightarrow 0$ and by Inequality 40 and the definition of $G^{*}$ we have